# Limits of Probability Measures with General Coefficients

### v2

@@ 1. Introduction
The goal of this paper is to characterize the convergence of a sequence of probability measures in terms of moments using the coefficients of their Bessel generating functions. In order to reach this goal, we study the applications of sequences of operators to formal power series. In particular, we consider the Dunkl operator introduced in [dunkloperators], which we define in Definition 3.1. Furthermore, we mainly focus on formal power series over $x_{i}$, $1\leq i\leq N$ that are symmetric in $N-1$ or $N$ of the $x_{i}$ for positive integers $N$. Partitions are useful for characterizing such formal power series, see Section 1.3.
@@ 1.1. Partitions
The set of all partitions is denoted by $P$. For $\lambda\in P$, $|\lambda|$ denotes the size of $\lambda$, and $P^{+}$ is defined as the set of $\lambda\in P$ such that $|\lambda|\geq 1$. We discuss some terminology and notation related to partitions that we use later.
@@ 1.2. Main result
Next, we state the setting of the main result. The setting we consider is based on the setting of the paper [matrix], and we state the definitions of LLN-satisfaction and exponentially decaying probability measures from the paper, see Definitions 1.2 and 1.3.
Suppose $\theta\in\mathbb{C}$ and $(\theta)\geq 0$. If $\beta=2\theta$, $\beta=1$, $2$, and $4$ correspond to the GUE, GOE, and GSE, respectively. For positive integers $N$, let $\mathcal{M}_{N}$ denote the set of Borel probability measures over $\mathbb{R}^{N}$. Given $\theta$ and $\mu\in\mathcal{M}_{N}$, the Bessel generating function $G_{\theta}(x_{1},\ldots,x_{N};\mu)$ is defined in Definition 4.3.
where $(a_{1},\ldots,a_{N})$ is distributed according to $\mu_{N}$ for $N,k\geq 1$. The moments $\{m_{k}\}_{k\geq 1}$ of $\{\mu_{N}\}_{N\geq 1}$ are given by $m_{k}=\lim_{N\rightarrow\infty}\mathbb{E}_{\mu_{N}}(p_{k}^{N,c})$ for $k\geq 1$.
Definition 1.2.
A sequence $\{\mu_{N}\}_{N\geq 1}$ of probability measures such that $\mu_{N}\in\mathcal{M}_{N}$ for $N\geq 1$ satisfies a Law of Large Numbers with constant $c$ if
(1) $\lim_{N\rightarrow\infty}\mathbb{E}_{\mu_{N}}\left(\prod_{i=1}^{s}p_{k_{i}}^{N,c}\right)=\prod_{i=1}^{s}m_{k_{i}}$
for any positive integer $s$ and positive integers $k_{i}$, $1\leq i\leq s$. For simplicity, we also equivalently state that the sequence $\{\mu_{N}\}_{N\geq 1}$ satisfies a $c$-LLN.
Definition 1.3.
A probability measure $\mu$ in $\mathcal{M}_{N}$ is exponentially decaying with exponent $R>0$ if
$\int_{\mathbb{R}^{N}}e^{NR\max_{1\leq i\leq N}|a_{i}|}\mu(da_{1},\ldots,da_{N})$
is finite.
Remark 1.4.
In addition to exponentially decaying probability measures, the results of this paper are also true for compactly supported generalized functions, see [rectangularmatrix]*Theorem 2.27.
Furthermore, we state Lemma 1.5, which is a key property of exponentially decaying probability measures. Afterwards, we state Theorem 1.7, which we use to prove Theorem 1.9, the main result of this paper. Note that in Theorem 1.7, $NC(k)$ denotes the set of noncrossing partitions of $[k]$, see Section 2.2 for the definition.
Lemma 1.5 ([matrix]*Lemma 2.9).
If $\mu\in\mathcal{M}_{N}$ is exponentially decaying with exponent $R>0$, then $G_{\theta}(x_{1},\ldots,x_{N};\mu)$ converges for all $(x_{1},\ldots,x_{N})$ in the domain
$\Omega_{R}=\left\{(x_{1},\ldots,x_{N})\in\mathbb{C}^{N}\big{|}|\text{Re}(x_{i})|<R,1\leq i\leq N\right\},$
and is holomorphic in the domain.
Remark 1.6.
For the proof of the lemma when $(\theta)\geq 0$, see Remark 4.5.
@@
Theorem 1.7 generalizes Claim 9.1 of the paper [matrix] and is similar to but also has important differences with Theorem 3.8 of the paper. Particularly, Theorem 3.8 and Claim 9.1 of the paper correspond to the cases $c=1$ and $c=0$, respectively. Furthermore, similarly to the proof of Theorem 3.8 of the paper, we prove Theorem 1.7 by evaluating Bessel generating functions at $x_{i}=0$, $1\leq i\leq N$ after applying Dunkl operators as well as other operators. Additionally, we consider when the limit of any partial derivative can be nonzero. On the other hand, in Theorem 3.8 and Claim 9.1 of the paper, the limits of partial derivatives with two or more distinct indices are 00. The arbitrary limits of partial derivatives are a reason why formal power series and in particular Theorem 3.2 are needed to prove Theorem 1.7.
Furthermore, [representations, fluctuations, unitary, llnclt, qmk, jackgenfunc, perelomov-popov, rectangularmatrix] consider similar results characterized by the limits of partial derivatives of Bessel, Jack, and Schur generating functions. If the partial derivatives involve only index, the limit can be nonzero, and in the papers [fluctuations, unitary, llnclt] studying the Central Limit Theorem, if two distinct indices are involved, the limit can be nonzero. However, if three or more distinct indices are involved, the limit must be zero. In this paper, we consider when the limits of partial derivatives involving any number of distinct indices can be nonzero.
In Theorem 1.7 the free cumulant of order $k\geq 1$ is
@@
using the definition of free cumulants presented in [freeprobability]. Observe that $c_{k}$ is a linear combination of the limits of the order $k$ partial derivatives. For the proof of Theorem 1.7, see Section 4.2.
In Section 4, we see that we can view $\ln(G_{\theta}(x_{1},\ldots,x_{N};\mu))$ as a symmetric polynomial and can therefore apply formal power series. Also, we often evaluate functions at $x_{i}=0$, $1\leq i\leq N$; for example, see (2). This corresponds to the constant terms of formal power series, and Theorem 3.2 is essential for this approach.
Furthermore, using Theorem 1.7 allows for the proof of the following generalization, see Section 6. The result requires an uniformity condition on the coefficients of the logarithms of the Bessel generating functions, similarly to what [llnclt]*Assumption 2.1 requires for Jack generating functions.
@@
Assume that for all $\nu\in P^{+}$, there exists $c_{\nu}\in\mathbb{R}_{\geq 0}$ such that $|c_{\nu}^{N}|\leq c_{\nu}+o_{N}(1)$. Then, $\{\mu_{N}\}_{N\geq 1}$ satisfies a $c$-LLN with free cumulants $\{c_{k}\}_{k\geq 1}$ if and only if
@@
for all $k\geq 1$.
For the previous result, recall that if the free cumulants are $\{c_{k}\}_{k\geq 1}$, then the moments are $m_{k}=\sum_{\pi\in NC(k)}\prod_{B\in\pi}c_{|B|}$ as well as the definition of $c$-LLN-satisfaction, see Definition 1.2.
@@ 1.3. Formal Power Series
The most important coefficient of a formal series $F(x_{1},\ldots,x_{N})$ in this paper is the constant term $[1]F(x_{1},\ldots,x_{N})=c_{F}^{(0,\ldots,0)}$. Particularly, we study the asymptotics of the constant terms of formal series after applying sequences of operators in the $N\rightarrow\infty$ limit.
We also consider formal series that are symmetric or symmetric in $N-1$ of the $x_{i}$. Suppose $N$ is a positive integer. For $\vec{x}=(x_{1},\ldots,x_{N})$ and $\nu\in P$ such that $\ell(\nu)\leq N$, let
@@
We focus on $F(x_{1},\ldots,x_{N})$ symmetric in $N-1$ of the $x_{i}$ in this paper. Moreover, with Definition 1.11, it is possible to consider the limits of sequences of formal series as $N$ increases to infinity. Particularly, the limit of a sequence $\{F_{N}(x_{1},\ldots,x_{N})\}_{N\geq i}$ with $F_{N}(x_{1},\ldots,x_{N})\in\mathcal{F}_{i}^{N}$ for $N\geq i$ can be viewed as $F_{i}(s)$ for some $s$ if the limit exists. This concept is discussed more in Section 2.
@@ 1.4. Paper Organization
The organization of this paper is as follows. In Section 2, we discuss results on applying sequences of operators to formal series with $N$ variables which are symmetric in $N-1$ variables as $N$ increases to $\infty$. Afterwards, in Section 3, we prove results involved in the proof of Theorem 1.7, and in Section 4, we prove the theorem. Following this, in Section 5, we apply Theorem 1.7 to the $\beta$-Hermite ensemble and discuss a generalization of the theorem, see Corollary 5.4. Finally, in Section 6, we consider the coefficients resulting from applications of Dunkl operators as polynomials in $\theta$ and $N$ and prove Theorem 1.9 as well as a generalization of the theorem, see Corollary 6.6.
@@ 2. Limits of Operators
In this section, we look at operators on formal series in $x_{j}$, $1\leq j\leq N$ such that there exists $i$, $1\leq i\leq N$ for which the formal series is symmetric for $j\not=i$. Recall that the set of such formal series is $\mathcal{F}_{i}^{N}$. The main results in this section are Theorem 2.9 and Theorem 2.13, with Theorem 2.9 being a special case of Theorem 2.13.
Furthermore, in this section, $\theta\in\mathbb{C}$ and the sequence $\{\theta_{N}\}_{N\geq 1}$ of complex numbers satisfies $\lim_{N\rightarrow\infty}\theta_{N}=\theta$. Note that $\theta$ does not need to have nonnegative real part, which is in contrast to the statements of Theorem 1.7 and Theorem 1.9.
@@ 2.1. Basic Results
For $N$ variables $x_{1},\ldots,x_{N}$, let $s_{i,j}$ be the operator which switches $x_{i}$ and $x_{j}$, for distinct $i,j\in[N]$. Note that $s_{i,j}$ is essential for the definition of the Dunkl operator, see Definition 3.1. Furthermore, for $1\leq i\leq N$, let $d_{i}$ be the operator such that for nonnegative integers $a_{k}$, $1\leq k\leq N$,
@@
In Proposition 2.3, we show that $\mathcal{Q}_{i}^{N}(f(x_{1},\ldots,x_{N}))$ is an operator from $\mathcal{F}_{i}^{N}$ to $\mathcal{F}_{i}^{N}$.
@@
Under certain conditions, the $\mathcal{Q}_{i}^{N}$ operators are asymptotically equivalent to Dunkl operators, see the proof of Theorem 3.17. We analyze them rather than Dunkl operators directly for simplicity. It would be interesting if results similar to this section’s results are true for Dunkl operators.
@@
Observe that $g_{2}(x_{1},\ldots,x_{N})$ is a formal power series which is symmetric in $\vec{x}_{i}$. We find the coefficient of $M_{\nu}(\vec{x}_{i})x_{i}^{d}$. If $\ell(\nu)\geq N$, the coefficient is 00 in $g_{2}(x_{1},\ldots,x_{N})$. Then, suppose $\ell(\nu)\leq N-1$. For a monomial $p$ in $M_{\nu}(\vec{x}_{i})$ with variables $(\vec{x}_{i})_{j}$, $1\leq j\leq\ell(\nu)$, there are $N-1-\ell(\nu)$ $j$ such that $x_{j}$ is not in $p$, and these are the $j$ such that $C_{i,j}$ applied to a monomial will give $px_{i}^{d}$. For a monomial $q$, if $C_{i,j}q=px_{i}^{d}$, then $q=px_{j}^{d+1}$ and has coefficient $c_{g}^{0,\nu+(d+1)}$ in $g$. Therefore, the coefficient of $px_{i}^{d}$, and thus $M_{\nu}(\vec{x}_{i})x_{i}^{d}$, in $g_{2}(x_{1},\ldots,x_{N})$ is $(N-\ell(\nu)-1)c_{g}^{0,\nu+(d+1)}$. From this,
@@
For $d\geq 0$ and $\nu\in P$, we find the coefficient of $M_{\nu}(\vec{x}_{i})x_{i}^{d}$ in $g_{3}(x_{1},\ldots,x_{N})$. If $\ell(\nu)\geq N$, the coefficient will be 00. Suppose $\ell(\nu)\leq N-1$, and let
@@
Let $m\geq 0$ be an integer. Suppose $\{f_{j,N}(x_{1},\ldots,x_{N})\}_{N\geq i}$ for $1\leq j\leq m$ and $\{g_{N}(x_{1},\ldots,x_{N})\}_{N\geq i}$ are symmetric outside of $i$. Then, $\ \left\{\left(\prod_{j=1}^{m}\mathcal{Q}_{i}^{N}(f_{j,N})\right)g_{N}(x_{1},\ldots,x_{N})\right\}_{N\geq i}$ is symmetric outside of $i$. Also, the limit outside of $i$ of $\mathcal{Q}_{i}^{N}(f_{1,N})g_{N}(x_{1},\ldots,x_{N})$ as $N\rightarrow\infty$ is
@@
The $m=0$ case is clear. For $m=1$, from Proposition 2.3, it is clear that the expression is symmetric outside of $i$. Furthermore, the result implies that the limit of $\mathcal{Q}_{i}^{N}(f_{1,N})g_{N}(x_{1},\ldots,x_{N})$ as $N\rightarrow\infty$ with respect to $i$ exists and equals (7). The rest of Corollary 2.4 can be proved using induction. ∎
@@
From Corollary 2.4, both limits exist. Suppose that for $1\leq j\leq m$, $f_{j}=\{c_{f_{j}}^{d,\nu}\}_{d\geq 0,\nu\in P}$, and $g=\{c_{g}^{d,\nu}\}_{d\geq 0,\nu\in P}$. We use induction on $m$, where the base case $m=0$ is clear. For $m=1$, we can use (7) to show that $\lim_{N\rightarrow\infty}\mathcal{Q}_{i}^{N}(f_{1,N})g_{N}(x_{1},\ldots,x_{N})=\lim_{N\rightarrow\infty}\mathcal{Q}_{i}^{N}(F(f_{1}))F(g)(x_{1},\ldots,x_{N})$.

### v3

@@ 1. Introduction
The goal of this paper is to characterize the convergence of a sequence of probability measures in terms of moments using the coefficients of their Bessel generating functions. In order to reach this goal, we study the applications of sequences of operators to formal power series. In particular, we consider the Dunkl operator introduced in [dunkloperators], which we define in Definition˜3.1. Furthermore, we mainly focus on formal power series over $x_{i}$, $1\leq i\leq N$ that are symmetric in N1N-1 or $N$ of the $x_{i}$ for positive integers $N$. Partitions are useful for characterizing such formal power series, see Section˜1.5.
1.1. Setup
We first state the setting of the main result. The setting we study is based on the setting of the paper [matrix], and we state the definitions of LLN-satisfaction and exponentially decaying measures from the paper in Definitions˜1.1 and 1.2, respectively, with some modifications; in particular, we prove results for a larger class of measures. The notion of LLN-satisfaction that we study is also discussed in [llnclt] in the context of Jack generating functions.
Suppose $\theta\in\mathbb{C}$ and $(\theta)\geq 0$. If $\beta=2\theta$, $\beta=1$, $2$, and $4$ correspond to the GUE, GOE, and GSE, respectively. For positive integers $N$, let $\mathcal{M}_{N}$ denote the set of Borel probability measures over $\mathbb{C}^{N}$. Given $\theta$ and $\mu\in\mathcal{M}_{N}$, the Bessel generating function $G_{\theta}(x_{1},\ldots,x_{N};\mu)$ is defined in Definition˜4.4.
where $(a_{1},\ldots,a_{N})\sim\mu_{N}$ for $N,k\geq 1$. The moments $\{m_{k}\}_{k\geq 1}$ of $\{\mu_{N}\}_{N\geq 1}$ are given by $m_{k}=\lim_{N\rightarrow\infty}\mathbb{E}[p_{k}^{N,c}]$ for $k\geq 1$.
Definition 1.1.
A sequence $\{\mu_{N}\}_{N\geq 1}$ of probability measures such that $\mu_{N}\in\mathcal{M}_{N}$ for $N\geq 1$ satisfies a Law of Large Numbers with constant $c$ and moments $\{m_{k}\}_{k\geq 1}$ if
$\lim_{N\rightarrow\infty}\mathbb{E}_{(a_{1},\ldots,a_{N})\sim\mu_{N}}\left(\prod_{i=1}^{s}p_{k_{i}}^{N,c}\right)=\prod_{i=1}^{s}m_{k_{i}}$
for all positive integers $s$ and $k_{i},1\leq i\leq s$. For simplicity, we also equivalently state that the sequence $\{\mu_{N}\}_{N\geq 1}$ satisfies a $c$-LLN.
For $v\in\mathbb{C}^{N}$, we let $|v|$ denotes its magnitude $\sqrt{\sum_{i=1}^{N}|v_{i}|^{2}}$. This notation is used in the following definition of the class of probability measures that we study.
Definition 1.2.
A probability measure $\mu$ in $\mathcal{M}_{N}$ is exponentially decaying at rate $R>0$ if
$\int_{\mathbb{C}^{N}}e^{R|(a_{1},\ldots,a_{N})|}\mu(da_{1},\ldots,da_{N})$
is finite.
Remark 1.3.
[dunklbound]*Lemma 4.4 discusses a condition similar to that of the previous definition. The results of this paper are also true for compactly supported generalized functions, see [rectangularmatrix]*Theorem 2.27.
The following lemma showcases a key property of exponentially decaying probability measures. It can be proved using the method of [matrix]*Lemma 2.9 and (18). Recall that we are assuming that $(\theta)\geq 0$.
Lemma 1.4.
Suppose $\mu\in\mathcal{M}_{N}$ is exponentially decaying at rate $R>0$. Then, $G_{\theta}(x_{1},\ldots,\\ x_{N};\mu)$ converges and is holomorphic in the closed ball of radius $R$ centered at the origin.
@@ 1.2. Partitions
Before stating the main result, we introduce some terminology and notation related to partitions. The set of all partitions is denoted by $P$. For $\lambda\in P$, $|\lambda|$ denotes the size of $\lambda$, and $P^{+}$ is defined as the set of $\lambda\in P$ such that $|\lambda|\geq 1$.
@@ 1.3. Main result
First, we state Theorem˜1.6, which we use to prove Theorem˜1.8, the main result of this paper. Note that in Theorem˜1.6, $NC(k)$ denotes the set of noncrossing partitions of $[k]$, see Section˜2.2 for the definition.
@@
In Theorem˜1.6 the free cumulant of order $k\geq 1$ is
@@
using the definition of free cumulants presented in [freeprobability]. Observe that $c_{k}$ is a linear combination of the limits of the order $k$ partial derivatives. For the proof of Theorem˜1.6, see Section˜4.2.
In Section˜4, we see that we can view $\ln(G_{\theta}(x_{1},\ldots,x_{N};\mu))$ as a symmetric polynomial and therefore as a symmetric formal power series. Also, we often evaluate functions at $x_{i}=0$, $1\leq i\leq N$; for example, see (1). This corresponds to the constant terms of formal power series, and Theorem˜3.2 is essential for this approach.
Using Theorem˜1.6 allows for the proof of the following generalization, see Section˜6. The result requires an uniformity condition on the coefficients of the logarithms of the Bessel generating functions, similarly to what [llnclt]*Assumption 2.1 requires for Jack generating functions.
@@
Assume that for all $\nu\in P^{+}$, $|c_{\nu}^{N}|=N^{o_{N}(1)}$. Then, $\{\mu_{N}\}_{N\geq 1}$ satisfies a $c$-LLN with free cumulants $\{c_{k}\}_{k\geq 1}$ if and only if
@@
for all $k\geq 1$. Recall that if the free cumulants are $\{c_{k}\}_{k\geq 1}$, then the moments are $m_{k}=\sum_{\pi\in NC(k)}\prod_{B\in\pi}c_{|B|}$ for $k\geq 1$.
In Section˜6, we also discuss a generalization of the previous theorem to the regime $|\theta_{N}N|\rightarrow\infty$, see Theorem˜6.13.
1.4. Related works
Theorem˜1.6 generalizes Claim 9.1 of the paper [matrix]. Furthermore, similarly to the proof of Theorem 3.8 of the paper, which corresponds to the case $c=1$, we prove Theorem˜1.6 by evaluating Bessel generating functions at $x_{i}=0$, $1\leq i\leq N$ after applying Dunkl operators as well as other operators. Additionally, we consider when the limit of any partial derivative can be nonzero. On the other hand, in Theorem 3.8 and Claim 9.1 of the paper, the limits of partial derivatives with two or more distinct indices are 0. The arbitrary limits of partial derivatives are a reason why formal power series and in particular Theorem˜3.2 are needed to prove Theorem˜1.6.
Additionally, [representations, fluctuations, unitary, llnclt, qmk, gaussianfluctuation, jackgenfunc, perelomov-popov, rectangularmatrix] consider similar results characterized by the limits of partial derivatives of Bessel, Jack, and Schur generating functions. If the partial derivatives involve only one index, the limit can be nonzero, and in the papers [fluctuations, unitary, llnclt, gaussianfluctuation] studying the Central Limit Theorem, if two distinct indices are involved, the limit can be nonzero. However, if three or more distinct indices are involved, the limit must be zero. In this paper, we consider when the limits of partial derivatives involving any number of distinct indices can be nonzero.
The paper [gaussianfluctuation] shows LLN and CLT results for the case $c=0$ and $\theta=1$ and does so after evaluating the Bessel generating function at a limiting distribution as $N\rightarrow\infty$ rather than at the origin. In particular, the paper evaluates the logarithm of the scaled Bessel generating function
$\mathbb{E}_{(a_{1},\ldots,a_{N})\sim\mu}\left[\frac{B_{(a_{1},\ldots,a_{N})}(x_{1},\ldots,x_{N};1)}{B_{(a_{1},\ldots,a_{N})}(\chi;1)}\right]$
at $(x_{1},\ldots,x_{N})=\chi$ for $\chi\in\mathbb{R}^{N}$; we consider the case when $\chi=(0,\ldots,0)$. However, we believe that the results of this paper are true for general values of $\chi$, since the results of Sections˜2 and 3 are true in this setting.
We introduce the operator $\mathcal{Q}_{i}^{N}$ in Section˜2 which computes the leading order terms after applying Dunkl operators to formal power series symmetric in all but one variable. The main contribution of this paper is the computation in the section of the leading order terms of the constant term of formal power series after the application of products of the $\mathcal{Q}_{i}^{N}$ operators.
@@ 1.5. Formal Power Series
The most important coefficient of a formal series $F(x_{1},\ldots,x_{N})$ in this paper is the constant term $1]F(x_{1},\ldots,x_{N})=c_{F}^{(0,\ldots,0)}$. Particularly, we study the asymptotics of the constant terms of formal series after applying sequences of operators in the $N\rightarrow\infty$ limit.
We also consider formal series that are symmetric or symmetric in N1N-1 of the $x_{i}$. Suppose $N$ is a positive integer. For $\vec{x}=(x_{1},\ldots,x_{N})$ and $\nu\in P$ such that $\ell(\nu)\leq N$, let
@@
We focus on $F(x_{1},\ldots,x_{N})$ symmetric in N1N-1 of the $x_{i}$ in this paper. Moreover, with Definition˜1.10, it is possible to consider the limits of sequences of formal series as $N$ increases to infinity. Particularly, the limit of a sequence $\{F_{N}(x_{1},\ldots,x_{N})\}_{N\geq i}$ with $F_{N}(x_{1},\ldots,x_{N})\in\mathcal{F}_{i}^{N}$ for $N\geq i$ can be viewed as $F_{i}(s)$ for some $s$ if the limit exists. This concept is discussed more in Section˜2.
@@ 1.6. Paper Organization
The organization of this paper is as follows. In Section˜2, we discuss results on applying sequences of operators to formal series with $N$ variables which are symmetric in N1N-1 variables as $N$ increases to $\infty$. Afterwards, in Section˜3, we prove results involved in the proof of Theorem˜1.6, and in Section˜4, we prove the theorem. Following this, in Section˜5, we apply Theorem˜1.6 to the $\beta$-Hermite ensemble and discuss a generalization of the theorem, see Corollary˜5.4. Finally, in Section˜6, we consider the coefficients resulting from applications of Dunkl operators as polynomials in $\theta$ and $N$ and prove Theorem˜1.8 and discuss a generalization to the regime $|N\theta_{N}|\rightarrow\infty$.
@@ 2. Limits of Operators
In this section, we look at operators on formal series in $x_{j}$, $1\leq j\leq N$ such that there exists $i$, $1\leq i\leq N$ for which the formal series is symmetric for $j\not=i$. Recall that the set of such formal series is $\mathcal{F}_{i}^{N}$. The main results in this section are Theorem˜2.9 and Theorem˜2.13, with Theorem˜2.9 being a special case of Theorem˜2.13.
Furthermore, in this section, $\theta\in\mathbb{C}$ and the sequence $\{\theta_{N}\}_{N\geq 1}$ of complex numbers satisfies $\lim_{N\rightarrow\infty}\theta_{N}=\theta$. Note that $\theta$ does not need to have nonnegative real part, which is in contrast to the statements of Theorem˜1.6 and Theorem˜1.8.
@@ 2.1. Basic Results
For $N$ variables $x_{1},\ldots,x_{N}$, let $s_{i,j}$ be the operator which switches $x_{i}$ and $x_{j}$, for distinct $i,j\in[N]$. Note that $s_{i,j}$ is essential for the definition of the Dunkl operator, see Definition˜3.1. Furthermore, for $1\leq i\leq N$, let $d_{i}$ be the operator such that for nonnegative integers $a_{k}$, $1\leq k\leq N$,
@@
In Proposition˜2.3, we show that $\mathcal{Q}_{i}^{N}(f(x_{1},\ldots,x_{N}))$ is an operator from $\mathcal{F}_{i}^{N}$ to $\mathcal{F}_{i}^{N}$.
@@
Under certain conditions, the $\mathcal{Q}_{i}^{N}$ operators are asymptotically equivalent to Dunkl operators, see the proof of Theorem˜3.17. We analyze them rather than Dunkl operators directly for simplicity. It would be interesting if results similar to this section’s results are true for Dunkl operators.
@@
Observe that $g_{2}(x_{1},\ldots,x_{N})$ is a formal power series which is symmetric in $\vec{x}_{i}$. We find the coefficient of $M_{\nu}(\vec{x}_{i})x_{i}^{d}$. If $\ell(\nu)\geq N$, the coefficient is 0 in $g_{2}(x_{1},\ldots,x_{N})$. Then, suppose $\ell(\nu)\leq N-1$. For a monomial $p$ in $M_{\nu}(\vec{x}_{i})$ with variables $(\vec{x}_{i})_{j}$, $1\leq j\leq\ell(\nu)$, there are $N-1-\ell(\nu)$ $j$ such that $x_{j}$ is not in $p$, and these are the $j$ such that $C_{i,j}$ applied to a monomial will give $px_{i}^{d}$. For a monomial $q$, if $C_{i,j}q=px_{i}^{d}$, then $q=px_{j}^{d+1}$ and has coefficient $c_{g}^{0,\nu+(d+1)}$ in $g$. Therefore, the coefficient of $px_{i}^{d}$, and thus $M_{\nu}(\vec{x}_{i})x_{i}^{d}$, in $g_{2}(x_{1},\ldots,x_{N})$ is $(N-\ell(\nu)-1)c_{g}^{0,\nu+(d+1)}$. From this,
@@
For $d\geq 0$ and $\nu\in P$, we find the coefficient of $M_{\nu}(\vec{x}_{i})x_{i}^{d}$ in $g_{3}(x_{1},\ldots,x_{N})$. If $\ell(\nu)\geq N$, the coefficient will be 0. Suppose $\ell(\nu)\leq N-1$, and let
@@
[…]
@@
The $m=0$ case is clear. For $m=1$, from Proposition˜2.3, it is clear that the expression is symmetric outside of $i$. Furthermore, the result implies that the limit of $\mathcal{Q}_{i}^{N}(f_{1,N})g_{N}(x_{1},\ldots,x_{N})$ as $N\rightarrow\infty$ with respect to $i$ exists and equals (6). The rest of Corollary˜2.4 can be proved using induction. ∎
@@
From Corollary˜2.4, both limits exist. Suppose that for $1\leq j\leq m$, $f_{j}=\{c_{f_{j}}^{d,\nu}\}_{d\geq 0,\nu\in P}$, and $g=\{c_{g}^{d,\nu}\}_{d\geq 0,\nu\in P}$. We use induction on $m$, where the base case $m=0$ is clear. For $m=1$, we can use (6) to show that $\lim_{N\rightarrow\infty}\mathcal{Q}_{i}^{N}(f_{1,N})g_{N}(x_{1},\ldots,x_{N})=\lim_{N\rightarrow\infty}\mathcal{Q}_{i}^{N}(F(f_{1}))F(g)(x_{1},\ldots,x_{N})$.